OpenStudy (anonymous):

a 4ft long rectangular table has a perimeter of 12ft the width of the rectangle is

OpenStudy (anonymous):

Perimeter = 2L+2W If L = 4, and P ( for perimeter ) = 12 12 = 2(4) + 2W 12 = 8 + 2W 4 = 2W ( we subtracted 8 from both sides ) 2 = W ( we divided both sides by 2 )